OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 3i

OpenStudy (jdoe0001):

so, what's its conjugate anyway?

OpenStudy (anonymous):

would it be 1-3i?

OpenStudy (jdoe0001):

yeap, so the product, will be \(\bf (1+3i)(1-3i)\\ \textit{keep in mind that } (a^2-b^2) = (a-b)(a+b)\)

OpenStudy (anonymous):

would 1-9i be correct? or is there more to it then that?

OpenStudy (jdoe0001):

... what's \(\bf \large i\) equals to again?

OpenStudy (anonymous):

square root of -1

OpenStudy (jdoe0001):

I see... so \(\bf (1+3i)(1-3i)\\ \textit{keep in mind that } (a^2-b^2) = (a-b)(a+b)\\ (1+3i)(1-3i) \implies 1^2-(3i)^2 \implies 1-3^2i^2 \\\quad \\ \implies 1-(3\times3)(\sqrt{-1}\times \sqrt{-1}) \implies 1-(9)(\sqrt{(-1)^2})\) what would that give you?

OpenStudy (jdoe0001):

\(\large 1-(9)\left(\sqrt{(-1)^2}\right)\) what do you think?

OpenStudy (anonymous):

10

OpenStudy (jdoe0001):

\(\bf 1-(9)\left(\sqrt{(-1)^2}\right) \implies 1-9(-1) \implies 10\)

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

(1+3i)(1-3i)=1-9(i^2)=1+9=10 conj=1-3i